OpenStudy (anonymous):

Solve for x: 3(x + 1) = −2(x − 1) + 6

OpenStudy (luigi0210):

Distribute first

HanAkoSolo (jamierox4ev3r):

thats what i was going to say Jesus... >.<

HanAkoSolo (jamierox4ev3r):

ahahaha whatever :P

OpenStudy (luigi0210):

@cgurley86 don't do their work fir them

OpenStudy (luigi0210):

*for

HanAkoSolo (jamierox4ev3r):

yeah @cgurley86 ... be sure to review this link http://openstudy.com/code-of-conduct

HanAkoSolo (jamierox4ev3r):

oh and welcome to OS....lolthanks @Luis_Rivera

OpenStudy (anonymous):

k

HanAkoSolo (jamierox4ev3r):

so @Tayloresc sorry to get sidetracked, but you distribute the numbers in the parentheses with the number out side of it

HanAkoSolo (jamierox4ev3r):



HanAkoSolo (jamierox4ev3r):

you follow me so far @Tayloresc ? based from this example, can you distribute the other side of the equation?

HanAkoSolo (jamierox4ev3r):

....

OpenStudy (anonymous):

Oh I'm sorry @Jamierox4ev3r I was on a different webpage yes I do

HanAkoSolo (jamierox4ev3r):

cool so can you distribute both of the entire equation for me? If you want, I can demonstrate for you

OpenStudy (anonymous):

demonstrate pleased then I can do it :)

HanAkoSolo (jamierox4ev3r):

cool i already gave an example above (see drawing)but a good rule of thumb is below: 2(2x+2) = 4x+4. Now you try

OpenStudy (anonymous):

2x-2?

HanAkoSolo (jamierox4ev3r):

ah, but do not forget the negatives: it would be -2x+2

HanAkoSolo (jamierox4ev3r):

if you put the two together, what do you get @Tayloresc ? Hint: 3x+3=...

OpenStudy (anonymous):

I'm so confused right now :( @Jamierox4ev3r

HanAkoSolo (jamierox4ev3r):

well, if you put the two distributed parts of the equation together, it is 3x+3=-2x+2

HanAkoSolo (jamierox4ev3r):

from there you should start being able to solve

HanAkoSolo (jamierox4ev3r):

3x+3=-2x+2 +2x +2x --------- 5x+3=2 -3 -3 ------- 5x=-1 /5 /5 ---- x= -0.2, or \(\large\frac{-1}{5}\)

OpenStudy (anonymous):

3(x + 1) = -2(x - 1) + 6 3x + 3 = -2x + 2 + 6 3x + 3 = -2x + 8 3x + 2x = 8 - 3 5x = 5 x = 1 check.. 3(1 + 1) = -2(1 - 1) + 6 3(2) = -2(0) + 6 6 = 6 (correct) x = 1

OpenStudy (anonymous):

jamierox....you forgot about the 6